Assume that Home and Foreign produce two goods, TVs and cars, and use the information below to answer the following questions.

No-Trade Equilibrium

Home

Foreign

wageTV= 12

wageC= ?

wage*TV = ?

wage*C= 6

aLTV = 0.5

aLC = ?

aLTV * = ?

aLC* = 1

PTV = ?

PC = 4

PTV = 3

PC = ?

a. What is the unit labor requirement for cars in Home?

b. What is the no-trade relative price of TVs in Home?

c. What is the unit labor requirement for TVs in Foreign?

d. What is the no-trade relative price of TVs in Foreign?

e. Suppose that the free trade equilibrium relative price of TVs, is PTV/ PC= 1. Which good will Home export?

f. Which good will Foreign export?
Answer #1

a) For finding the unit labour requirement, we need to look at the equation

Pc= aLC* wage C

Wages in both sectors in home would be same because If there is difference in wages , workers would shift to the one with higher wages

Wage C= wageTv=12

4= aLC* 12

So aLC=1/ 3

Hence unit labour requirement is1/ 3 for cars at home.

b) The same equation which was used to find aLC would be used here to find Ptv

Ptv= aLTV*wageTv

Ptv= 0.5* 12

Ptv= 6

No trade relative Price of Tv is PTv/ Pc= 6/4= 3/2

c) formula for unit wage requirement in foreign is given by

Ptv*= wagetv* × aLTv*

3= 6× altv* ( wages in both sectors in foreign would be same)

aLTV*= 1/2

So unit labour requirement for TVs in foreign is1/2

d) Pc*= wageC*× alC*

Pc*= 6×1=6

So no trade relative price of tv in foreign is

PTv*/PC* = 3/6=1/2

e) We have the free trade relative price as PTv/Pc= 1

Sp the country having less of opportunity cost than the relative price will export.

In home market

alTV/aLC= 0.5/1/3=3/2

It is given that PTv/Pc=1

So Ptv/Pc < .aLT/aLC

aLTV*/ aLc*= 1/2/1=1/2

Ptv/PC> aLTv*/aLc*

The opportunity cost of producing Tv in home is more and of producing cloth is less.

Similarly the opportunity cost of producing Tv is less in foreign. So home produces and exports cloth.

f) From the same working above we know that foreign has a lower opportunity in producing Tv and hence would produce and export Tv.

(You can comment for doubts)